Nonconstant FCF Growth Valuation

Reizenstein Technologies (RT) has just developed a solar panel capable of generating 200% more electricity than any solar panel currently on the market. As a result, RT is expected to experience a 16% annual growth rate for the next 5 years. By the end of 5 years, other firms will have developed comparable technology, and RT's growth rate will slow to 7% per year indefinitely. RT has a 13% weighted average cost of capital. The most recent annual free cash flow (FCF0) was $3.20 million.

  1. Calculate RT's expected FCFs for t = 1, t = 2, t = 3, t = 4, and t = 5. Do not round intermediate calculations. Enter your answer in millions. For example, an answer of $1 million should be entered as 1, not 1,000,000. Round your answers to three decimal places.

    FCF1 = $    million

    FCF2 = $    million

    FCF3 = $    million

    FCF4 = $    million

    FCF5 = $    million

  2. What is the horizon value at t = 5 (HV5)? Do not round intermediate calculations. Enter your answer in millions. For example, an answer of $1 million should be entered as 1, not 1,000,000. Round your answers to two decimal places.

    $    million

  3. What is the present value of the horizon value (HV5)? Do not round intermediate calculations. Enter your answer in millions. For example, an answer of $1 million should be entered as 1, not 1,000,000. Round your answers to two decimal places.

    $    million

  4. What is the present value of the free cash flows expected at t = 1, t = 2, t = 3, t = 4,and t = 5? Do not round intermediate calculations. Enter your answer in millions. For example, an answer of $1 million should be entered as 1, not 1,000,000. Round your answers to two decimal places.

    $    million

  5. What is the value of operations at t = 0? Do not round intermediate calculations. Enter your answer in millions. For example, an answer of $1 million should be entered as 1, not 1,000,000. Round your answers to two decimal places.

    $    million
Answer #1

a.FCF 1 = FCF 0 (1+Growth rate)

= 3.20 million (1+16%) = $3.712 million

FCF2 = 3.20(1.16)2 = $4.306 million

FCF 3 = 3.20(1.16)3 = $4.995 million

FCF 4 = 3.20(1.16)4 = $5.794 million

FCF 5 = 3.20(1.16)5 = $6.721 million

b.Horizon value = 6.721(1+7%)/(13%-7%)

= $119.86 million

c.Present value = 119.86/(1.13)5

= $65.06 million

d.Present value = 3.712/(1.13) + 4.306/(1.13)2 + 4.995/(1.13)3 + 5.794/(1.13)4 + 6.721/(1.13)5

= $17.32 million

e.Value of operations = 65.06+17.32

= $82.38 millions